# Data-Driven Adaptive PID Control Based on Physics-Informed Neural Networks

### v1

@@ Abstract
{IEEEkeywords}
Physics-informed neural networks, Adaptive control, PID control, Model predictive control, Machine learning
1 Introduction
[…]
@@ 1 Introduction
The rest of the paper is structured as follows. Section 2 reviews PINNs for control [30] and its MPC-based implementation [26]. In Section 3, we propose an adaptive PID control led by the optimization problem with a PINNs-based dynamical model and show the implementation of the proposed PID control. In Sections 4 and 5, we demonstrate the effectiveness of the proposed control performance through simulations with two typical case studies of the 2-DOF manipulator systems and the MSD systems. Section 6 concludes the paper.
2 Preliminaries
2.1 Original Control Problem
@@
where $J$ is the stage cost and $\Psi$ is the terminal cost. Assume that the stage cost $J$ satisfies $\nabla_{\bm{u}^{2}}J>0$ uniformly at $\bm{u}\in\mathbb{U}$. The details of $J$ and $\Psi$ will be introduced in Section 3. The model-based optimal control problem (P1) with the nominal model-based state transition (2) can be solved numerically or partially-analytically (e.g., [32, 33]). However, it is challenging to incorporate the actual dynamics, which include unmodeled dynamics and parametric uncertainty, into the optimal control problem (P1).
2.2 Physics-Informed Neural Networks for Control
@@
Let us define the two training datasets. One is the time-series measured dataset $\mathcal{D}_{\text{data}}=\{(t_{i},\bm{x_{0,i}},\bm{x_{f,i}},\bm{u_{i}})\}$, $i\in\mathcal{L}_{\text{data}}=\{1,\ldots,N_{\text{data}}\}$, where $\bm{x_{0,i}}\in\mathbb{X}$ is the initial state included in the admissible state set $\mathbb{X}$, $\bm{x_{f,i}}$ is the final state at time $t_{i}\in(0,\Delta T+\epsilon]$, $\epsilon>0$, and $\bm{u_{i}}$ is the control input for the $i$-th data point. The other is the dataset $\mathcal{D}_{\text{phys}}=\{(t_{j},\bm{x_{j}},\bm{u_{j}})\}$, $j\in\mathcal{L}_{\text{phys}}=\{1,\ldots,N_{\text{phys}}\}$ uniformly and randomly self-selected over $0,\Delta T+\epsilon]\times\mathbb{X}\times\mathbb{U}$ to evaluate the physics-based dynamical model. Note that the PINNs model is trained for a period of length $\Delta T+\epsilon$, which is longer than the sampling time $\Delta T$, to ensure the automatic differentiation at the sampling time $\Delta T$ used in the optimization problem, which will be mentioned in Section 3. Then, the loss function for training the parameters $\omega$ in the PINNs model $\hat{\varphi}$ from the collected training dataset $\mathcal{D}_{\text{data}}$ and $\mathcal{D}_{\text{phys}}$ is defined as follows:
@@
3 PINNs-based Adaptive PID Control
@@
3.1 PID Gain Optimization Problem
@@
From the discussion in Section 2, the state trajectory (2) can be replaced with a PINNs-based prediction model $\hat{\varphi}$ in (6) and (9). Then, to stabilize the nonlinear system (1) over a wide state range, a time-varying PID gain optimization of the controller (11) given as the solution of the optimization problem (P1) is provided by the continuous-time formulation
@@
Appropriate selection of the admissible gain set $\mathcal{F}$ yields a controller that guarantees the internal stability of the systems. Compared to the nominal model (1), a PINNs-based model calibrated to real physical data can accommodate unmodeled dynamics and uncertainties more effectively. Consequently, the proposed approach is efficient when integrated with real-world data and is promising from a Sim2Real perspective. Meanwhile, it is challenging to obtain a PINNs-based, theoretically guaranteed, stabilized controller. A partial discussion regarding the guarantee of internal stability will be discussed in Section 5.3. Validation on real-data applications is left for future work.
3.2 Implementation
In this subsection, we introduce the methodology to implement the proposed adaptive PID control (P6). The typical system models (1) used for validation will be discussed in Sections 5 and 6.
@@ 3.2 Implementation
To implement the above methodology through simulations discussed in Sections 5 and 6, we utilize the virtual machine computer Proxmox Virtual Environment (Proxmox VE), which is recognized as one of the leading open-source virtualization platforms. [^1] The detailed information of Proxmox VE can be found at https://www.proxmox.com/en/products/proxmox-virtual- environment/overview As for the hardware environment, we allocated 10 cores from the host machine’s Intel Core i9-12900 processor (24 cores) to the virtual machine, along with 48GB of memory from a total of 128GB, and configured it to utilize the NVIDIA RTX 3090 through GPU passthrough directly. For the software environment, we conducted experiments on Linux and adopted TensorFlow 2.12.0 as the deep learning framework. All code was written in Python. To ensure reproducibility, all random number seed values are fixed at a common value.
Under the above settings, we evaluate the effectiveness of the proposed control methodology through simulations of two system models in Sections 4 and 5.
@@
4 Case Study 1: 2-DOF Robot Manipulator
@@
4.1 Properties on steady-state PID control
@@
4.2 Evaluation on PINNs-based model
@@
4.3 Application to Regulation Problem
@@
The time evolutions of the control input $\bm{u}(t)=[u_{\alpha}(t),u_{\beta}(t)]^{\top}$ and the corresponding proposed PID gains $\bm{F}_{k}$ are shown in Fig. 5 and Fig. 5, respectively. The fixed PID controller exhibits relatively low-frequency behavior during the transient state. In other words, the convergence to the steady-state of $\alpha$ is relatively poor because the initial response saturates at the upper input constraint. Conversely, the proposed method improves this situation by increasing the gain of the control input $u_{\beta}$ according to the size of the deviation and by finely adjusting the input within the constraint range. From the time evolutions of the proposed PID gains shown in Fig. 5, we observe different characteristics for transient and steady-state conditions. In the transient state, the proportional and differential gains are dominant, whereas in the steady state, the integral gain plays a significant role in stabilizing the system. The numerical results of the steady-state are equivalent to the tendency of the theoretical analysis mentioned in Section 4.1. The phenomenon of a temporary increase in $\bm{K}^{p}_{k}$ and $\bm{K}^{d}_{k}$ during the transient response can be interpreted as the result of dynamically optimizing the trade-off between the convergence property and the stabilization property.
@@
4.4 Application to Servo Problem
Finally, we evaluate the performance for the servo problem, which involves the tracking control problem with sequentially stepwise changes in target values. The reference trajectory $\bm{x}^{\text{ref}}=[\alpha^{\text{ref}},\beta^{\text{ref}},0,0]^{\top}$ is defined as a series of step changes in joint angles, as shown in Fig 6. The reference trajectory of the angles $\alpha^{\text{ref}},\beta^{\text{ref}}]^{\top}$ rad transitions to different target angles at specific time intervals; $1.0,0.0]^{\top}$ at $t=8$ s, $1.0,0.0]^{\top}$ at $16$ s, $0.5,0.0]^{\top}$ at $24$ s and $0,0]^{\top}$ at $32$ s. We emphasize that the reference trajectory is significantly affected by gravity and cannot be stabilized without control input. The other experimental conditions and the fixed PID method are the same as those for the regulation problem discussed in Section 4.3.
@@ 4.4 Application to Servo Problem
We next discuss the control input shown in Fig. 6 and the corresponding PID gain shown in Fig. 6. We see from the figures that the gains of the proposed PID method are automatically and high-frequently adjusted based on the online state information of the nonlinear system and the corresponding error between the real-time state and the stepwise reference trajectory. The numerical results of the dominant reference-dependent integral gain in steady state are equivalent to the tendency observed in the theoretical analysis mentioned in Section 4.1. Note that, as the proposed method can effectively handle the uncertainties and variations in the system dynamics by leveraging the data-driven nature of the PINNs-based prediction model and computing the optimization quickly using only data and automatic differentiation, the proposed method is decisively different from the conventional model-based approach.
@@
5 Case Study 2: Mass-Spring-Damper System
@@
5.1 Evaluation on PINNs-based model
The training dataset to obtain the model (6) is generated by the same method shown in Sections 3.2 and 4.2. We consider the optimization problem (P6) with the prediction horizon $T=4.0$ based on the recurrent self-loop PINNs prediction of the sampling period $\Delta T=0.200$ s. The training datasets, $\mathcal{D}_{\text{data}}$ and $\mathcal{D}_{\text{phys}}$, are composed of $N_{\text{data}}=2\times 10^{4}$ and $N_{\text{phys}}=1\times 10^{5}$ samples, respectively. These samples are generated with $\Delta T+\epsilon=0.250$ s.
@@ 5.1 Evaluation on PINNs-based model
5.2 Application to Servo Problem
We next evaluate the performance of the proposed method for the servo problem of the dynamical system (22), similar to Section 4.4. Note that in this servo problem, while the PINNs model was trained without noise consideration, we introduce 3% multiplicative Gaussian sensor noise in the dynamical system (22) to test the robustness of the control performance under measurement uncertainty. The reference trajectory $\bm{x}^{\text{ref}}=[z^{\text{ref}},0]^{\top}$ is defined as a series of step changes in position, as shown in Fig. 9. The position reference trajectory $z^{\text{ref}}$ transitions to different target positions at specific time intervals: $0.3$ at $t=20$ s, 0.5-0.5 at $40$ s. These transitions test the system’s ability to track both positive and negative step changes in the presence of measurement noise. We set the initial state $\bm{x}(0)=[-0.7,0.0]^{\top}$, and the control sets $\mathbb{U}=[-1.0,1.0]$, $\mathcal{F}=[0.0,5.0]\times[0.0,5.0]\times[0.0,5.0]$. The adaptive PID control is implemented during $60.0$ s using the MPC-based adaptive PID gain optimization (P6) with $\Delta T=0.200$ s, $t_{f}=60.0$ s, and the weight matrices of (13) given by $Q={\rm diag}(1000,1)$ and $R=0.01$. The optimizer to solve the problem (P6) uses Adam [34] with a learning rate of $10^{-2}$, and the maximum iteration is set as $2.0\times 10^{4}$. For comparison, we employ a conventional fixed-gain PID controller with $\bm{F}_{k}\equiv[K^{p}_{k},K^{i}_{k},K^{d}_{k}]=[1.2,1.0,1.2]$ for all time steps $k$. These gains are tuned empirically to achieve acceptable tracking performance while maintaining stability under the input constraints.
@@ 5.3 Stability Analysis
To address the stability issues identified above, we reformulate the servo problem from Section 5.2 by incorporating explicit stability-guaranteed constraints. Applying the Routh-Hurwitz stability criterion to the open-loop transfer function (23), the closed-loop stability conditions for the PID gains are derived as
@@
Under these circumstances, the resulting control performance, led by the reformulated PID gain optimization problem (P6) with the modified term (25), is shown in Figs. 12 and 13. The time evolution of the stability margin in Fig. 13 reveals that the obtained controller satisfies the stability-guaranteed constraint (24). As shown in Fig. 12, the stability-constrained adaptive PID controller demonstrates superior tracking performance compared to the fixed-gain PID controller, consistent with the results in Section 5.2. Moreover, when compared with the unconstrained adaptive case shown in Fig. 9, the stability-guaranteed formulation achieves better tracking accuracy while maintaining guaranteed closed-loop stability. From Fig. 12, the proportional and derivative gains maintain larger values throughout the entire time domain compared to the unconstrained case. The stability-guaranteed gains increase proportional and derivative action, resulting in an average gain crossover frequency of $3.27$ rad/s, approximately $2.8$ times higher than in the unconstrained case. The higher gain crossover frequency indicates increased closed-loop bandwidth, while the larger derivative gain introduces phase lead compensation, both contributing to the improved tracking performance observed in Fig. 12. For non-origin reference positions, the integral gain increases to eliminate steady-state errors, similar to the unconstrained approach, but within stability bounds. Fig. 13 presents the stability analysis for the stability-constrained case. The Nyquist plots in Figs. $13$ demonstrate that the open-loop frequency response maintains a safe distance from the critical point 1-1 throughout the operation, in contrast to the unconstrained case shown in Fig. 10. Fig. 13 confirms that the stability margin remains close to 1.0 for almost the entire time domain, confirming that the optimization successfully maintains sufficient stability margins while achieving the control objectives. To verify the robustness of the stability-constrained controller, we conducted the same disturbance rejection test as in Fig. 11. As shown in Fig. 14, the system maintains stability even under the constant disturbance, in stark contrast to the divergent behavior observed in the unconstrained case.
@@
These results suggest that properly constraining the PID gain search space to ensure stability can significantly enhance the performance of optimization-based gain design. While the Routh-Hurwitz criterion provides analytical stability-guaranteed constraints for linear systems, such as the MSD model, deriving these constraints for general nonlinear systems remains challenging, as demonstrated by the manipulator dynamics in Section 4, where analytical stability conditions are typically intractable. This limitation presents a significant challenge that must be addressed for the broader application of the proposed method to complex nonlinear systems, particularly when explicitly considering stability.
6 Conclusion

### v2

@@ Abstract
I Introduction
@@
The rest of the paper is structured as follows. Section II reviews PINNs for control [30] and its MPC-based implementation [26]. In Section III, we propose an adaptive PID control led by the optimization problem with a PINNs-based dynamical model and show the implementation of the proposed PID control. In Sections IV and V, we demonstrate the effectiveness of the proposed control performance through simulations with two typical case studies of the 2-DOF manipulator systems and the MSD systems. Section VI concludes the paper.
II Preliminaries
II-A Original Control Problem
@@
where $J$ is the stage cost and $\Psi$ is the terminal cost. Assume that the stage cost $J$ satisfies $\nabla_{\bm{u}^{2}}J>0$ uniformly at $\bm{u}\in\mathbb{U}$. The details of $J$ and $\Psi$ will be introduced in Section III. The model-based optimal control problem (P1) with the nominal model-based state transition (2) can be solved numerically or partially-analytically (e.g., [32, 33]). However, it is challenging to incorporate the actual dynamics, which include unmodeled dynamics and parametric uncertainty, into the optimal control problem (P1).
II-B Physics-Informed Neural Networks for Control
@@
Let us define the two training datasets. One is the time-series measured dataset $\mathcal{D}_{\text{data}}=\{(t_{i},\bm{x_{0,i}},\bm{x_{f,i}},\bm{u_{i}})\}$, $i\in\mathcal{L}_{\text{data}}=\{1,\ldots,N_{\text{data}}\}$, where $\bm{x_{0,i}}\in\mathbb{X}$ is the initial state included in the admissible state set $\mathbb{X}$, $\bm{x_{f,i}}$ is the final state at time $t_{i}\in(0,\Delta T+\epsilon]$, $\epsilon>0$, and $\bm{u_{i}}$ is the control input for the $i$-th data point. The other is the dataset $\mathcal{D}_{\text{phys}}=\{(t_{j},\bm{x_{j}},\bm{u_{j}})\}$, $j\in\mathcal{L}_{\text{phys}}=\{1,\ldots,N_{\text{phys}}\}$ uniformly and randomly self-selected over $0,\Delta T+\epsilon]\times\mathbb{X}\times\mathbb{U}$ to evaluate the physics-based dynamical model. Note that the PINNs model is trained for a period of length $\Delta T+\epsilon$, which is longer than the sampling time $\Delta T$, to ensure the automatic differentiation at the sampling time $\Delta T$ used in the optimization problem, which will be mentioned in Section III. Then, the loss function for training the parameters $\omega$ in the PINNs model $\hat{\varphi}$ from the collected training dataset $\mathcal{D}_{\text{data}}$ and $\mathcal{D}_{\text{phys}}$ is defined as follows:
@@
III PINNs-based Adaptive PID Control
@@
III-A PID Gain Optimization Problem
@@
From the discussion in Section II, the state trajectory (2) can be replaced with a PINNs-based prediction model $\hat{\varphi}$ in (6) and (9). Then, to stabilize the nonlinear system (1) over a wide state range, a time-varying PID gain optimization of the controller (11) given as the solution of the optimization problem (P1) is provided by the continuous-time formulation
@@
Appropriate selection of the admissible gain set $\mathcal{F}$ yields a controller that guarantees the internal stability of the systems. Compared to the nominal model (1), a PINNs-based model calibrated to real physical data can accommodate unmodeled dynamics and uncertainties more effectively. Consequently, the proposed approach is efficient when integrated with real-world data and is promising from a Sim2Real perspective. Meanwhile, it is challenging to obtain a PINNs-based, theoretically guaranteed, stabilized controller. A partial discussion regarding the guarantee of internal stability will be discussed in Section V-C. Validation on real-data applications is left for future work.
III-B Implementation
In this subsection, we introduce the methodology to implement the proposed adaptive PID control (P6). The typical system models (1) used for validation will be discussed in Sections V and VI.
@@ III-B Implementation
To implement the above methodology through simulations discussed in Sections V and VI, we utilize the virtual machine computer Proxmox Virtual Environment (Proxmox VE), which is recognized as one of the leading open-source virtualization platforms. [^1] The detailed information of Proxmox VE can be found at https://www.proxmox.com/en/products/proxmox-virtual- environment/overview As for the hardware environment, we allocated 10 cores from the host machine’s Intel Core i9-12900 processor (24 cores) to the virtual machine, along with 48GB of memory from a total of 128GB, and configured it to utilize the NVIDIA RTX 3090 through GPU passthrough directly. For the software environment, we conducted experiments on Linux and adopted TensorFlow 2.12.0 as the deep learning framework. All code was written in Python. To ensure reproducibility, all random number seed values are fixed at a common value.
Under the above settings, we evaluate the effectiveness of the proposed control methodology through simulations of two system models in Sections IV and V.
@@
IV Case Study 1: 2-DOF Robot Manipulator
@@
IV-A Properties on steady-state PID control
@@
IV-B Evaluation on PINNs-based model
@@
IV-C Application to Regulation Problem
@@
The time evolutions of the control input $\bm{u}(t)=[u_{\alpha}(t),u_{\beta}(t)]^{\top}$ and the corresponding proposed PID gains $\bm{F}_{k}$ are shown in Fig. 5 and Fig. 5, respectively. The fixed PID controller exhibits relatively low-frequency behavior during the transient state. In other words, the convergence to the steady-state of $\alpha$ is relatively poor because the initial response saturates at the upper input constraint. Conversely, the proposed method improves this situation by increasing the gain of the control input $u_{\beta}$ according to the size of the deviation and by finely adjusting the input within the constraint range. From the time evolutions of the proposed PID gains shown in Fig. 5, we observe different characteristics for transient and steady-state conditions. In the transient state, the proportional and differential gains are dominant, whereas in the steady state, the integral gain plays a significant role in stabilizing the system. The numerical results of the steady-state are equivalent to the tendency of the theoretical analysis mentioned in Section IV-A. The phenomenon of a temporary increase in $\bm{K}^{p}_{k}$ and $\bm{K}^{d}_{k}$ during the transient response can be interpreted as the result of dynamically optimizing the trade-off between the convergence property and the stabilization property.
@@
IV-D Application to Servo Problem
Finally, we evaluate the performance for the servo problem, which involves the tracking control problem with sequentially stepwise changes in target values. The reference trajectory $\bm{x}^{\text{ref}}=[\alpha^{\text{ref}},\beta^{\text{ref}},0,0]^{\top}$ is defined as a series of step changes in joint angles, as shown in Fig 6. The reference trajectory of the angles $\alpha^{\text{ref}},\beta^{\text{ref}}]^{\top}$ rad transitions to different target angles at specific time intervals; $1.0,0.0]^{\top}$ at $t=8$ s, $1.0,0.0]^{\top}$ at $16$ s, $0.5,0.0]^{\top}$ at $24$ s and $0,0]^{\top}$ at $32$ s. We emphasize that the reference trajectory is significantly affected by gravity and cannot be stabilized without control input. The other experimental conditions and the fixed PID method are the same as those for the regulation problem discussed in Section IV-C.
@@ IV-D Application to Servo Problem
We next discuss the control input shown in Fig. 6 and the corresponding PID gain shown in Fig. 6. We see from the figures that the gains of the proposed PID method are automatically and high-frequently adjusted based on the online state information of the nonlinear system and the corresponding error between the real-time state and the stepwise reference trajectory. The numerical results of the dominant reference-dependent integral gain in steady state are equivalent to the tendency observed in the theoretical analysis mentioned in Section IV-A. Note that, as the proposed method can effectively handle the uncertainties and variations in the system dynamics by leveraging the data-driven nature of the PINNs-based prediction model and computing the optimization quickly using only data and automatic differentiation, the proposed method is decisively different from the conventional model-based approach.
@@
V Case Study 2: Mass-Spring-Damper System
@@
V-A Evaluation on PINNs-based model
The training dataset to obtain the model (6) is generated by the same method shown in Sections III-B and IV-B. We consider the optimization problem (P6) with the prediction horizon $T=4.0$ based on the recurrent self-loop PINNs prediction of the sampling period $\Delta T=0.200$ s. The training datasets, $\mathcal{D}_{\text{data}}$ and $\mathcal{D}_{\text{phys}}$, are composed of $N_{\text{data}}=2\times 10^{4}$ and $N_{\text{phys}}=1\times 10^{5}$ samples, respectively. These samples are generated with $\Delta T+\epsilon=0.250$ s.
@@ V-A Evaluation on PINNs-based model
V-B Application to Servo Problem
We next evaluate the performance of the proposed method for the servo problem of the dynamical system (22), similar to Section IV-D.
Note that in this servo problem, while the PINNs model was trained without noise consideration, we introduce 3% multiplicative Gaussian sensor noise in the dynamical system (22) to test the robustness of the control performance under measurement uncertainty. The reference trajectory $\bm{x}^{\text{ref}}=[z^{\text{ref}},0]^{\top}$ is defined as a series of step changes in position, as shown in Fig. 9. The position reference trajectory $z^{\text{ref}}$ transitions to different target positions at specific time intervals: $0.3$ at $t=20$ s, 0.5-0.5 at $40$ s. These transitions test the system’s ability to track both positive and negative step changes in the presence of measurement noise. We set the initial state $\bm{x}(0)=[-0.7,0.0]^{\top}$, and the control sets $\mathbb{U}=[-1.0,1.0]$, $\mathcal{F}=[0.0,5.0]\times[0.0,5.0]\times[0.0,5.0]$. The adaptive PID control is implemented during $60.0$ s using the MPC-based adaptive PID gain optimization (P6) with $\Delta T=0.200$ s, $t_{f}=60.0$ s, and the weight matrices of (13) given by $Q={\rm diag}(1000,1)$ and $R=0.01$. The optimizer to solve the problem (P6) uses Adam [34] with a learning rate of $10^{-2}$, and the maximum iteration is set as $2.0\times 10^{4}$. For comparison, we employ a conventional fixed-gain PID controller with $\bm{F}_{k}\equiv[K^{p}_{k},K^{i}_{k},K^{d}_{k}]=[1.2,1.0,1.2]$ for all time steps $k$. These gains are tuned empirically to achieve acceptable tracking performance while maintaining stability under the input constraints.
@@ V-C Stability Analysis
To address the stability issues identified above, we reformulate the servo problem from Section V-B by incorporating explicit stability-guaranteed constraints. Applying the Routh-Hurwitz stability criterion to the open-loop transfer function (23), the closed-loop stability conditions for the PID gains are derived as
@@
Under these circumstances, the resulting control performance, led by the reformulated PID gain optimization problem (P6) with the modified term (25), is shown in Figs. 12 and 13. The time evolution of the stability margin in Fig. 13 reveals that the obtained controller satisfies the stability-guaranteed constraint (24). As shown in Fig. 12, the stability-constrained adaptive PID controller demonstrates superior tracking performance compared to the fixed-gain PID controller, consistent with the results in Section V-B. Moreover, when compared with the unconstrained adaptive case shown in Fig. 9, the stability-guaranteed formulation achieves better tracking accuracy while maintaining guaranteed closed-loop stability. From Fig. 12, the proportional and derivative gains maintain larger values throughout the entire time domain compared to the unconstrained case. The stability-guaranteed gains increase proportional and derivative action, resulting in an average gain crossover frequency of $3.27$ rad/s, approximately $2.8$ times higher than in the unconstrained case. The higher gain crossover frequency indicates increased closed-loop bandwidth, while the larger derivative gain introduces phase lead compensation, both contributing to the improved tracking performance observed in Fig. 12. For non-origin reference positions, the integral gain increases to eliminate steady-state errors, similar to the unconstrained approach, but within stability bounds. Fig. 13 presents the stability analysis for the stability-constrained case. The Nyquist plots in Figs. $13$ demonstrate that the open-loop frequency response maintains a safe distance from the critical point 1-1 throughout the operation, in contrast to the unconstrained case shown in Fig. 10. Fig. 13 confirms that the stability margin remains close to 1.0 for almost the entire time domain, confirming that the optimization successfully maintains sufficient stability margins while achieving the control objectives. To verify the robustness of the stability-constrained controller, we conducted the same disturbance rejection test as in Fig. 11. As shown in Fig. 14, the system maintains stability even under the constant disturbance, in stark contrast to the divergent behavior observed in the unconstrained case.
@@
These results suggest that properly constraining the PID gain search space to ensure stability can significantly enhance the performance of optimization-based gain design. While the Routh-Hurwitz criterion provides analytical stability-guaranteed constraints for linear systems, such as the MSD model, deriving these constraints for general nonlinear systems remains challenging, as demonstrated by the manipulator dynamics in Section IV, where analytical stability conditions are typically intractable. This limitation presents a significant challenge that must be addressed for the broader application of the proposed method to complex nonlinear systems, particularly when explicitly considering stability.
VI Conclusion